OpenStudy (luckycharms988):

PLEASE HELP. Solve for x. x + 2/ 3 = -4

OpenStudy (anonymous):

Hey!

OpenStudy (luckycharms988):

Oh hey lol.

OpenStudy (anonymous):

2/3 is being added to 'x' What's the opposite of addition? We need to do the opposite to isolate 'x'

OpenStudy (luckycharms988):

subtract x and 2/3?

OpenStudy (anonymous):

Partially correct, we only subtract 2/3 x + 2/3 - 2/3 = x -4 - 2/3 = ??

OpenStudy (anonymous):

x = -4 - 2/3

OpenStudy (luckycharms988):

-3?

OpenStudy (anonymous):

Nope, here: -4 - 2/3 = -1( 4 + 2/3)

OpenStudy (luckycharms988):

oh, -12?

OpenStudy (luckycharms988):

i mean -14

OpenStudy (anonymous):

no, let's make 4 have the same denominator 4 = 12/3 12/3 + 2/3 = ?

OpenStudy (luckycharms988):

14/3 and simplified would be 4 2/3

OpenStudy (anonymous):

and then we have to multiply that negative 1 that we had factored out so our final answer is -14/3 or -4 2/3

OpenStudy (luckycharms988):

okay thanks so much

OpenStudy (anonymous):

You're welcome! :)